OpenStudy (anonymous):

Using differential calculus; calculate the time to reach the greatest voltage rate of change when v=100*sin(200*pi*t+pi/4)

OpenStudy (anonymous):

The voltage rate of change is just dv/dt This will have max, min, inflection points when d^2v / dt^2 = 0 or is undefined.

OpenStudy (anonymous):

Thanks for your help, So when d^2v/dt^2 = 4*10^6*pi^2*sin(200*pi*t+pi/4) = 0, I solve for t and that should give me the time at which the greatest voltage rate of change occurs?